An experienced mechanic earns $8.00 per hour more than an apprentice. If it takes the apprentice 18 hours to earn the same amount of money as the mechanic can earn in 12 hours, what is the hourly pay rate of the mechanic?
Answer by ankor@dixie-net.com(6693)

An experienced mechanic earns $8.00 per hour more than an apprentice.
If it takes the apprentice 18 hours to earn the same amount of money as
the mechanic can earn in 12 hours, what is the hourly pay rate of the mechanic?
:
Let x = hourly amt earned by an experienced mechanic
then
(x-8) = amt earned by an apprentice
:
Mech amt earned in 12 hr = appren amt earned in 18 hrs
12x = 18(x-8)
12x = 18x - 144
144 = 18x - 12x
144 = 6x
x =
x = $24 an hour
;
:
Check solution
12(24) = 18(16)